# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v1

@@ Abstract
The realization of universal robots is an ultimate goal of researchers. However, a key hurdle in achieving this goal lies in the robots’ ability to manipulate objects in their unstructured surrounding environments according to different tasks. The learning-based approach is considered an effective way to address generalization. The impressive performance of foundation models in the fields of computer vision and natural language suggests the potential of embedding foundation models into manipulation tasks as a viable path toward achieving general manipulation capability. However, we believe achieving general manipulation capability requires an overarching framework akin to auto driving. This framework should encompass multiple functional modules, with different foundation models assuming distinct roles in facilitating general manipulation capability. This survey focuses on the contributions of foundation models to robot learning for manipulation. We propose a comprehensive framework and detail how foundation models can address challenges in each module of the framework. What’s more, we examine current approaches, outline challenges, suggest future research directions, and identify potential risks associated with integrating foundation models into this domain.
@@ 1 Introduction
Researchers aim to create universal robots that can seamlessly integrate into human life to boost productivity, much like those depicted in the movie ”I, Robot”. However, a key hurdle in achieving this lies in the robots’ ability to manipulate objects in their unstructured surrounding environments according to different tasks. There is abundant literature available for improving the general manipulation capabilities of robots, which can be roughly categorized into model-based and learning-based approaches (Zarrin et al. (2023)). The real world is too diverse for universal robots and they must adapt to unstructured environments and arbitrary objects to manipulate effectively. Therefore, learning-based methods are crucial for manipulation tasks (Kleeberger et al. (2020)).
The predominant methodologies in learning-based approaches are deep learning, reinforcement learning and imitation learning. Learning-based methods have spanned from acquiring specific manipulation skills through labeled datasets like human demonstration, to acquiring abstract representations of manipulation tasks conducive to high-level planning, to exploring an object’s functionalities through interaction and encompassing various objectives in between (Kroemer et al. (2021)). However, challenges persist, including 1) unnatural interaction with humans; 2) high-cost data collection; 3) limited perceptual capabilities; 4) non-intelligent hierarchy of skills; 5) inaccurate pre- and post-conditions & post-hoc correction; 6) unreliable skill learning; 7) poor environment transition (Hu et al. (2023b)).
@@
    Large Language Models (LLMs) like BERT (Devlin et al. (2018)), GPT-3 (Brown et al. (2020)) have demonstrated the capability to generate coherent chains of thought.
@@
    Vision Foundation Models (VFMs) like SAM (Kirillov et al. (2023)) demonstrate strong segmentation capabilities for open-set objects.
@@
    Robot Foundation Models (RFMs), like RT-X (Padalkar et al. (2023a)). Internet-scale dataset, such as images and text data, are suitable for pre-training visual and language models, but lack task level manipulation data. Therefore, researchers aim to train end-to-end RFMs by collecting task-level manipulation datasets to enable observation-to-action mapping.
@@
    RFMs serve as an end-to-end policy model, directly outputting actions based on input observations
@@
These findings underscore the potential of embedding foundation models into manipulation tasks as a viable path toward achieving general manipulation capabilities. However, we do not believe that a single foundation model alone can achieve general manipulation capabilities. Although RFMs currently represent a single-model end-to-end training approach, ensuring safety and stability, particularly in achieving an over 99% safety rate in manipulation tasks, remains a challenge. Therefore, drawing inspiration from the development of autonomous driving systems (Hu et al. (2023c)), achieving general manipulation capabilities necessitates an overarching framework that encompasses multiple functional modules, with different foundation models assuming distinct roles in facilitating general manipulation capabilities.
The ultimate general manipulation framework should be capable of interacting with humans to manipulate arbitrary objects in open-world scenarios and achieve diverse manipulation objectives. However, the interaction between agent and human involves not only recognizing intentions but also learning new skills or improving old skills from human experts in the external world. Open-world scenarios can be static or dynamic, with objects being either rigid or deformable and task objectives ranging from short-horizon to long-horizon. We designate the restriction of the agent’s learning capability to improving old skills and manipulating rigid objects in static scenes to achieve short-horizon task objectives as Level 0 (L0) . Drawing from Kroemer et al. (2021), we propose a comprehensive framework for general manipulation task at L0. Hence, we aim to use this survey not only to enlighten scholars on the issues that foundation models can address in robot learning for manipulation but also to stimulate their exploration of the general manipulation framework and the role various foundation models can play in the general manipulation framework.
@@
This paper is structured as follows: In Sec. 2, we present a comprehensive framework of robot learning for general manipulation, based on the developmental history of robot learning. We elaborate on the impact of foundation models on each module in the framework in the following sections. Sec. 3 is Human/Agnet Interaction module, Sec. 4 is Pre- and Post-conditions Detection module, Sec. 5 is Hierarchy of Skills module, Sec. 6 is State Perception module, Sec. 7 is Policy module, Sec. 8 is Manipulation Data Generation module. In Sec. 9, we discuss several issues of particular concern to us. In Sec. 10, we summarize the contributions of this survey and identify the limitations of the current framework as well as the ongoing challenges.
@@ 2 Framework of Robot Learning for General Manipulation
Over the past decade, there has been a significant expansion in research concerning robot manipulation, with a focus on leveraging the growing accessibility of cost-effective robot arms and grippers to enable robots to interact directly with the environment in pursuit of their objectives. As the real world encompasses extensive variation, a robot cannot expect to possess an accurate model of its surrounding environment, the objects within it, or the skills necessary for manipulation in advance (Kroemer et al. (2021)).
@@
Early stage, robotic manipulation is defined as learning a policy $\Pi$ through deep learning, reinforcement learning, or imitation learning, etc. This policy controls the robot’s joint movements and executes tasks based on observations of the environment and the robot’s state $S$, mapping to actions $\alpha$. such as Rlafford (Geng et al. (2023)) and Graspnet (Fang et al. (2020)) take point cloud as input and output the target poses.
@@
In the mid-term, many tasks in robotics require a series of correct actions, which are often long-horizon tasks. For example, making a cup of tea with a robot involves multiple sequential steps such as boiling water, adding a tea bag, pouring hot water, etc. Learning to plan for long horizons is a central challenge in episodic learning problems (Wang et al. (2020)). Decomposing tasks has several advantages. It makes learning individual skills more efficient by breaking them into shorter-horizon, thus aiding exploration. Reusing skills in multiple settings can speed up learning by avoiding the need to relearn elements from scratch each time. Researchers train a hierarchy model $H$ to decompose the task into a sequence of subgoals (Ahn et al. (2022)), and observe pre- and post-conditions to ensure that the prerequisites and outcomes of each subgoals are met (Cui et al. (2022)). This comes from human experience in task completion, breaking tasks down into several steps and being able to decide on the execution method for each step based on the initial environment, such as putting a tea bag into a teacup using either a pick-place or pushing method (pre-condition), and ultimately detecting whether the step is completed to proceed to the next step (post-condition).
@@
Recently, researchers have realized that training policies require real-world interaction between robots and environments, which inevitably increases the probability of unforeseen hazardous situations. Therefore, researchers aim to train the environment’s transition model $T$. Once the model is fitted, agents can generate samples based on it, significantly reducing the frequency of direct interaction between agents and environments (Liu et al. (2024)).
@@
We organize the framework of robot learning for general manipulation according to its development history, as shown in Fig. 2. Due to the task instruction and corrective instruction are commanded by humans or other agents, we include an interaction module $I$ in the framework. The aforementioned modules all require datasets for learning. The data collection process for the hierarchy of skills $H$ and pre- and post-conditions detection modules $P$ is similar to that in the fields of CV and NLP. Compared to data collection in CV and NLP domains, gathering datasets for manipulation tasks involves interaction between robots and the real environment, leading to high costs and risks. A simulation environment can mitigate the challenges of dataset collection for manipulation tasks, but there still exists a sim-to-real gap. Therefore, we emphasize the manipulation data generation $D$, to provide dataset for the state $S$, policy $P$, and transition $T$ modules.
@@
    Human/Agent Interaction. As a general-purpose robot, it is necessary to understand the task goal instruction conveyed by humans or collaborative agents to accomplish tasks. It is also important to synchronize and receive corrective instructions when one’s actions deviate from the goal or when the goal changes during task execution.
@@
    Pre- and Post-conditions Detection. Before the hierarchy of skills and executing a skill, check the environment to ensure it meets the requirements and select the execution method based on object affordances. For instance, when placing a tea bag in a teacup, pre-conditions detection determines whether there are tea bags and teacups and chooses between pick-place or pushing based on their affordances, such as, a tea bag is spherical, and it has the affordance of rolling when pushed. After executing a skill, verify if the objective is achieved. If not, post-conditions detection identifies the reason for failure. For example, if a teacup is knocked over during pick-and-place, inform post-hoc correction and use pick-and-place to upright the cup and reinsert the tea bag.
@@
    Hierarchy of skills. The concept of ’Hierarchy of skills’ often involves creating a sequence of subgoals (Song et al. (2023)). Each subgoal necessitates a skill, which may consist of one or multiple primitive actions (Zhang et al. (2023c)). For instance, tasks like filling the kettle with water, heating the water, and getting the tea leaves are examples of subgoals that agents need to achieve in a specific order to fulfill the final goal as instructed.
@@
    State. The state consists of robot proprioception $S_{robot}$, general environmental state $S_{e}$, and object states $S_{o}$. The difference between $S_{e}$ and $S_{o}$ is analogous to the foreground and background of an image. $S_{robot}$ generally relates to the mechanical structure of robots. Currently, there are limited studies focusing on the improvement of robot mechanical structures using foundation models, with Stella et al. (2023) being one of them. However, researches in this direction are scarce and still in their initial stages.
@@
    Policy. The policy outputs actions to accomplish task goals based on the input state. At presents, the paradigms for generating actions using foundation models include LLMs directly generating code for robot execution, VLMs directly generating or VLMs combined with LLMs generating corresponding target poses, RFMs directly outputting target pose or delta pose through end-to-end training, and foundation models assisting reinforcement learning in generating corresponding actions.
@@
    Action. Different foundation models applied as policies result in different types of actions. We categorize them into three types: Code, Target Pose, and Delta Pose. Code refers to the direct control code of the robot arm and gripper. Target Pose refers to the desired pose of the end-effector, which is input to motion planning to generate the trajectory. Delta Pose refers to the next waypoint the end-effector moves to, with continuously outputted delta poses forming the trajectory.
@@
    Transition. The transition model can directly output the changed state of the environment and objects based on the input action, thus helping to reduce the interaction between the robot and the real environment. Zhang and Soh (2023) uses LLM as a human model, predicting the changes in human state after robot actions, and selecting the most suitable robot actions for interaction with humans. In various algorithms, the transition environment model is not a mandatory module. Therefore, in Fig. 2, we enclose this module with a dashed box.
@@
    Manipulation Data Generation. The manipulation data generation module provides manipulation of task-level datasets.
Figure 2: Framework of Robot Learning for General Manipulation. Human or other agents convey task instruction or corrective instruction to the Hierarchy of Skills module $H$ through the Interaction module $I$. Pre-conditions Detection module $P$ conveys object detection and affordance detection perception information to Hierarchy of Skills $H$. Hierarchy of Skills may encounter ambiguity in generating subgoals, such as when there are two cups in the environment, it needs to determine which cup to pour water into through interaction with the external world. After generating subgoals, Policy module $\Pi$ trains according to the dataset generated by Manipulation Data Generation module $D$. During Skill Execution, Action $\alpha$ is generated by perceiving State $S$. Environment Transition $T$ is not a necessary module; its addition can be directly transitioned to the next state through the generated action. When issues arise during execution, Corrective Instruction is conveyed to Policy module $\Pi$ from the external world. Policy module $\Pi$ modifies the current action to corrective action and saves corrective demonstration to the dataset for self-improvement of Policy module $\Pi$. After skill execution, Post-conditions Detection module $P$ determines the success of execution. If successful, proceed to the next step; if not, the failure reason is conveyed to Post-hoc Correction module for self-correction.

### v2

@@ Abstract
The realization of universal robots is an ultimate goal of researchers. However, a key hurdle in achieving this goal lies in the robots’ ability to manipulate objects in their unstructured environments according to different tasks. The learning-based approach is considered an effective way to address generalization. The impressive performance of foundation models in the fields of computer vision and natural language suggests the potential of embedding foundation models into manipulation tasks as a viable path toward achieving general manipulation capability. However, we believe achieving general manipulation capability requires an overarching framework akin to auto driving. This framework should encompass multiple functional modules, with different foundation models assuming distinct roles in facilitating general manipulation capability. This survey focuses on the contributions of foundation models to robot learning for manipulation. We propose a comprehensive framework and detail how foundation models can address challenges in each module of the framework. What’s more, we examine current approaches, outline challenges, suggest future research directions, and identify potential risks associated with integrating foundation models into this domain.
@@ 1 Introduction
Researchers aim to create universal robots that can seamlessly integrate into human life to boost productivity, much like those depicted in the movie ”I, Robot”. However, a key hurdle in achieving this lies in the robots’ ability to manipulate objects in their unstructured environments according to different tasks. There is abundant literature available for improving the general manipulation capability of robots, which can be roughly categorized into model-based and learning-based approaches (Zarrin et al. (2023)). The real world is too diverse for universal robots and they must adapt to unstructured environments and arbitrary objects to manipulate effectively. Therefore, learning-based methods are crucial for manipulation tasks (Kleeberger et al. (2020)).
The predominant methodologies in learning-based approaches are deep learning, reinforcement learning and imitation learning. Learning-based methods have spanned from acquiring specific manipulation skills through labeled datasets like human demonstration, to acquiring abstract representations of manipulation tasks conducive to high-level planning, to exploring an object’s functionalities through interaction and encompassing various objectives in between (Kroemer et al. (2021)). However, challenges persist, including 1) unnatural interaction with humans; 2) high-cost data collection; 3) limited perceptual capability; 4) non-intelligent hierarchy of skills; 5) inaccurate pre- and post-conditions & post-hoc correction; 6) unreliable skill learning; 7) poor environment transition (Hu et al. (2023b)).
@@
    Large Language Models (LLMs) like BERT (Devlin et al. (2018)), GPT-3 (Brown et al. (2020)) demonstrate the capability to generate coherent chains of thought.
@@
    Vision Foundation Models (VFMs) like SAM (Kirillov et al. (2023)) demonstrate strong segmentation capability for open-set objects.
@@
    Robot Foundation Models (RFMs), like RT-X (Padalkar et al. (2023a)). Internet-scale dataset, such as images and text data, are suitable for pre-training visual and language models, but lack task level manipulation data. Therefore, researchers aim to train end-to-end RFMs by collecting task-level manipulation datasets to enable observations-to-action mapping.
@@
    RFMs serve as an end-to-end policy model, directly outputting actions based on input blue observations.
@@
These findings underscore the potential of embedding foundation models into manipulation tasks as a viable path toward achieving general manipulation capability. However, we do not believe that a single foundation model alone can achieve general manipulation capability. Although RFMs currently represent a single-model end-to-end training approach, ensuring safety and stability, particularly in achieving an over 99% success rate in manipulation tasks, remains a challenge. Achieving over a 99% success rate in manipulation tasks is crucial, as human manipulation success rates are around 99%. Without this level of accuracy, robots can’t replace humans. Therefore, drawing inspiration from the development of autonomous driving systems (Hu et al. (2023c)), achieving general manipulation capability necessitates an overarching framework that encompasses multiple functional modules, with different foundation models assuming distinct roles in facilitating general manipulation capability.
The ultimate general manipulation framework should be able to interact with human or other agent and to manipulate arbitrary objects in open-world scenarios and achieve diverse manipulation tasks. However, the interaction between robot and human involves not only recognizing intentions but also learning new skills or improving old skills from human experts in the external world. Open-world scenarios may be static or dynamic. Objects can be either rigid or deformable. Task objectives can vary from short-term to long-term. Furthermore, tasks may necessitate different degrees of precision with respect to contact points and applied forces/torques. We designate the restriction of the robot’s learning capability to improving old skills and to manipulating rigid objects in static scenes in order to achieve short-horizon task objectives with low precision requirements for contact points and forces/torques as Level 0 (L0). Drawing from Kroemer et al. (2021), we propose a comprehensive framework for general manipulation task at L0. Hence, we aim to use this survey not only to enlighten scholars on the issues that foundation models can address in robot learning for manipulation but also to stimulate their exploration of the general manipulation framework and the role various foundation models can play in the general manipulation framework.
@@
This paper is structured as follows: In Sec. 2, we present a comprehensive framework of robot learning for general manipulation, based on the developmental history of robot learning for manipulation and general manipulation definition. We elaborate on the impact of foundation models on each module in the framework in the following sections. Sec. 3 is Human/Agnet Interaction module, Sec. 4 is Pre- and Post-conditions Detection module, Sec. 5 is Hierarchy of Skills module, Sec. 6 is State Perception module, Sec. 7 is Policy module, Sec. 8 is Manipulation Data Generation module. In Sec. 9, we discuss several issues of particular concern to us. In Sec. 10, we summarize the contributions of this survey and identify the limitations of the current framework as well as the challenges in each module.
@@ 2 Framework of Robot Learning for General Manipulation
Over the past decade, there has been a significant expansion in research concerning robot manipulation, with a focus on leveraging the growing accessibility of cost-effective robot arms and grippers to enable robots to interact directly with the environment in pursuit of their objectives. As the real world encompasses extensive variation, a robot cannot expect to possess an accurate model of its unstructured environment, the objects within it, or the skills necessary for manipulation in advance (Kroemer et al. (2021)).
@@
Early stage, robot manipulation is defined as learning a policy $\Pi$ through deep learning, reinforcement learning, or imitation learning, etc. This policy controls the robot’s joint movements and executes tasks based on observations of the environment and the robot’s state $S$, mapping to actions $\alpha$. such as Rlafford (Geng et al. (2023)) and Graspnet (Fang et al. (2020)) take point cloud as input and output the target pose. This process is represented by the Skill Execution module, as shown in Fig. 2.
@@
In the mid-term, many tasks in robotics require a series of correct actions, which are often long-horizon tasks. For example, making a cup of tea with a robot involves multiple sequential steps such as boiling water, adding a tea bag, pouring hot water, etc. Learning to plan for long-horizon tasks is a central challenge in episodic learning problems (Wang et al. (2020)). Decomposing tasks has several advantages. It makes learning individual skills more efficient by breaking them into shorter-horizon, thus aiding exploration. Reusing skills in multiple settings can speed up learning by avoiding the need to relearn elements from scratch each time. Researchers train a hierarchy model to decompose the task into a sequence of subgoals (Ahn et al. (2022)), and observe pre- and post-conditions to ensure that the prerequisites and outcomes of each subgoals are met (Cui et al. (2022)). These three processes are represented as the Hierarchy of Skills module $H$, the Pre-conditions Detection module $P$, and the Post-conditions Detection module $P$ in Fig. 2. However, detecting only task success with post-conditions detection is insufficient. It should also identify the reasons for task failure to help the robot self-correct and improve success rates. Therefore, we add a Post-hoc Correction module, as shown in Fig. 2.
@@
Recently, researchers have realized that training policies require real-world interaction between the robot and environments, which inevitably increases the probability of unforeseen hazardous situations. Therefore, researchers aim to train the environment’s transition model $T$. Once the model is fitted, robot can generate samples based on it, significantly reducing the frequency of direct interaction between the robot and environments (Liu et al. (2024)). This process is represented as the Transition module $T$ in Fig. 2.
The modules described above are summarized from the development of robot learning for manipulation. However, they are still insufficient for a comprehensive framework for general manipulation. The ultimate general manipulation framework should be able to interact with human or other agent and control whole-body to manipulate arbitrary objects in open-world scenarios, achieving diverse manipulation tasks. When interacting with human or other agent to understand task objectives, the transmitted instruction may sometimes be unclear, such as when there are two cups in the environment, it needs to determine which cup to pour water. Therefore, we add the Interaction module $I$ in Fig. 2 to understand the precise task objective.
@@
The aforementioned modules all require datasets for learning. The data collection process for the Hierarchy of skills $H$ and Pre- and Post-conditions detection modules $P$ is similar to that in the fields of CV and NLP. Compared to data collection in CV and NLP domains, gathering datasets for manipulation tasks requires the robot’s trajectory to train the policy. Therefore, we include the Manipulation Data Generation module in Fig. 2.
We organize the framework of robot learning for general manipulation according to its development history and definition, as shown in Fig. 2. In the caption of Fig. 2, we outline the flow of the entire framework. To better illustrate the role of each module, we list the inputs and outputs of each module below, along with their specific functions.
@@
    Pre-conditions Detection. This module takes raw information observed by the robot as input. It outputs perception information about objects in the environment and affordances of those object. Perception information helps ensures that requirements are met and helps select the execution method based on object affordances. For instance, when placing a tea bag in a teacup, perception information can help determine whether there are tea bags and teacups and chooses between pick-place or pushing based on their affordances, such as, a tea bag is spherical, and it has the affordance of rolling when pushed.
@@
    Human/Agent Interaction. The input to the Human/Agent Interaction module $I$ consists of an instruction or answer from the collaborating agent or human, and perception information from the Pre-conditions Detection module $P$. The output includes a question if the instruction or answer has ambiguities and provides a precise instruction to the Hierarchy of Skills module $H$. The main function of this module is to understand the exact task objectives.
@@
    Hierarchy of skills. This module takes as input the perception information about objects in the environment and their affordances for the task from the Pre-conditions Detection module $P$, as well as the precise instruction from the Interaction module $I$. It then produces a sequence of subgoals as output. The concept of ’Hierarchy of skills’ often involves creating a sequence of subgoals (Song et al. (2023)). Each subgoal necessitates a skill, which may consist of one or multiple primitive actions (Zhang et al. (2023c)). For instance, tasks like filling the kettle with water, heating the water, and getting the tea leaves are examples of subgoals that robot needs to achieve in a specific order to fulfill the final goal as instructed.
@@
    State. The input to the State module is the current environment, objects and robot states. States require the use of multiple sensors for perception. The output is the features of states. The states consists of robot proprioception $S_{robot}$, environment state $S_{e}$, and objects states $S_{o}$. The difference between $S_{e}$ and $S_{o}$ is analogous to the foreground and background of an image. $S_{robot}$ generally relates to the mechanical structure of the robot. Currently, there are limited studies focusing on the improvement of robot mechanical structure using foundation models, with Stella et al. (2023) being one of them. However, researches in this direction are scarce and still in their initial stages.
@@
    Policy. The Policy module takes as input features from the State module $S$ and subgoals generated by the Hierarchy of Skills module $H$. The policy outputs action to accomplish task goals based on the input states. We categorize action into three types: Code, Target Pose, and Delta Pose. Code refers to the direct control code of the robot. Target Pose refers to the desired pose of the end-effector, which is input to motion planning to generate the trajectory. Delta Pose refers to the next waypoint the end-effector moves to, with continuously outputted delta pose forming the trajectory. At present, the methods for generating actions using foundation models include LLMs directly generating code for robot execution, VLMs directly generating or VLMs combined with LLMs generating corresponding target poses, RFMs directly outputting target pose or delta pose through end-to-end training, and foundation models assisting reinforcement learning in generating various actions.
@@
    Post-conditions Detection. This module takes as input the environment, objects and robot states observed after the robot performs a task, along with the subgoals generated by the Hierarchy of Skills module H. It outputs whether the current subgoal is successful. If not, it provides the reason for failure to Post-hoc Correction module. The Post-hoc Correction module generates a sequence of actions for self-correction based on the failure reason. For example, if a teacup is knocked over during pick-and-place, inform post-hoc correction and use pick-and-place to upright the cup and reinsert the tea bag.
@@
    Transition. The Transition module $T$ takes an action generated by the Policy module $P$ as input. It outputs the next state after executing this action, thus helping to reduce the interaction between the robot and the real environment. Zhang and Soh (2023) uses LLM as a human model, predicting the changes in human state after robot actions, and selecting the most suitable robot actions for interaction with humans. In various algorithms, the transition environment model is not a mandatory module. Therefore, in Fig. 2, we enclose this module with a dashed box.
@@
    Manipulation Data Generation. This module functions as a database. It takes in existing manipulation data and correction data generated from robot tasks. The output is to provide task-level manipulation datasets for offline training.
Current research on foundation models for manipulation primarily focuses on several key modules: the Interaction module in Sec. 3, the Pre- and Post-conditions Detection module in Sec. 4, the Hierarchy of Skills module in Sec. 5, the State module in Sec. 6, the Policy module in Sec.7, and the Manipulation Data Generation module in Sec. 8. The following section will provide an overview of these modules.
Figure 2: Framework of Robot Learning for General Manipulation. The Pre-conditions Detection module $P$ perceives the environment to identify objects and the affordances objects support. The Interaction module $I$ receives instruction from a human or other agent. It uses perception information from the Pre-conditions Detection module $P$ to check for ambiguities in the instruction. If there are any ambiguities, it generates a question to clarify the instruction by asking the human or other agent. The Hierarchy of Skills module $H$ generates subgoals by using precise instruction from the Interaction module $I$ and perception information from the Pre-conditions Detection module $P$. Each subgoal is then passed to the Skill Execution module. In the Skill Execution module, Policy module $\Pi$ generates Action $\alpha$ based on the State $S$. To obtain the next state after executing the current action, State $S$ can either perceive it from the environment or use the Transition module $T$. To train the Skill Execution module, including the State module $S$, the Policy module $\Pi$ and the Transition module $T$, the Manipulation Data Generation module is required. This module provides a task-level manipulation dataset. When issues arise during execution, corrective instruction is sent to the Policy module $\Pi$ for manual adjustment. Policy module $\Pi$ modifies the current action to corrective action and saves corrective demonstration to the dataset for self-improvement of Policy module $\Pi$. After skill execution, Post-conditions Detection module $P$ determines the success of execution. If successful, proceed to the next subgoal; if not, the failure reason is conveyed to Post-hoc Correction module for self-correction.